OpenStudy (anonymous):

L1:(-1,6),(1,4) and L2:(3,-3),(6,-9) are the plots on the lines parallel, perpendicular, or neither?

OpenStudy (anonymous):

not the plots but the lines and they give the plots to find out if they lines are one of those!

OpenStudy (anonymous):

Label each point of L1: x1=-1 y1=6, x2=1 y=4 Compute m1 the slope of L1. m1=deltay/deltax deltay=y2-y1 deltaX=x2-x1 deltay=(4)-(6)=-2 deltax=1-(-1)=1+1=2 deltay/deltax=-2/2=-1 So m1 = -1 Now, if m2 the slope of L2=-1 (also) the the lines are parallel if m2=+1 the line are perpendicular If m2 is neither then the lines are neither parallel nor perpendicular. m2=(-9)-(-3) / (6)-(3) m2= -9+3/6-3 m2=-6/3=-2. Conclusion, the lines are neither parallel nor perpendicular. Lots more problem like this solved in math911 www.math911.com